# Interpretable Embeddings of Speech Enhance and Explain Brain Encoding Performance of Audio Models

### v1

@@ 4 Discussion
Recent papers (Anderson, Davis, and Lalor 2024; Goldstein et al. 2025) have claimed a paradigm shift from context-invariant handcrafted features to SSM embeddings in building computational models of speech perception. However, our findings caution against treating SSM embeddings as a standalone complete model for speech perception. We showed that interpretable models alone can outperform SSMs in predicting brain responses, especially in brain regions responsible for processing low-level acoustic information. Our results showed that the combination of SSM embeddings and interpretable features achieved the best predictive performance, highlighting the importance of the continued use of interpretable features in building computational models of human speech perception.
One of our main contributions is to explicitly demonstrate that SSMs preferentially preserve low-level acoustic information within the speech critical bands (100 – 1000 Hz). Previous studies (Pasad, Chou, and Livescu 2022; Pasad, Shi, and Livescu 2023) on the layer-wise representation in SSMs analyzed acoustic information more broadly, without examining frequency-specific selectivity. To the best of our knowledge, the closest related findings are: Wang et al. (Wang and Yamagishi 2022) showed that speech anti-spoofing systems based on SSMs were relatively robust to band-stop filtering of high frequencies but not to filtering in the 0-0.8 kHz range. Our results build on these observations by providing direct evidence of frequency-selective representations within SSMs, showing that the preservation of low-frequency components is not only critical for performance but also intrinsically encoded in the model’s internal representations.
@@
Whether SSMs encode brain-relevant semantics is debated, with evidence supporting (Pasad, Chou, and Livescu 2022; Chen et al. 2024; Anderson, Davis, and Lalor 2024) and challenging (Choi et al. 2024; Oota et al. 2024). Our method employs an interpretable semantic embedding to directly test for the shared neural variance with SSMs, while controlling for lower-level acoustic and linguistic information. We show that SSM representations significantly reduce the unique variance explained by the semantic embedding, providing strong evidence that they indeed encode brain-relevant semantic information. However, an important nuance remains: SSMs account for only about 40% of the unique variance explained by the semantic features, indicating that their semantic representations remain partial. Recent studies (Moussa, Klakow, and Toneva 2025; Vattikonda et al. 2025) have shown that fine-tuning SSMs on fMRI data can improve their alignment with language regions of the brain. Thus, while our findings affirm the presence of semantic information in SSM representations, they also suggest room for improvement.
Limitations
@@
1. Our findings are based on data from nine clinical ECoG participants who listened to a single 30‑min English podcast. While the high temporal resolution of ECoG and broad electrode coverage across brain regions make this dataset particularly well-suited for the present study, replication using larger and more diverse datasets will be important for assessing the generalizability of these results. 2. While we controlled for several lower-level features in our analysis of semantic encoding in SSMs, we cannot entirely rule out the possibility that unaccounted confounds contribute to the shared variance between SSM representations and our semantic features. However, the observed effect of context size provides some evidence against this interpretation, as lower-level features typically operate over shorter integration windows and are therefore less sensitive to extended context. 3. While our interpretable features account for most of the neural variance captured by SSM representations, a statistically significant portion remains unexplained. Identifying the nature of this residual variance is an important direction for future research.
Acknowledgements
The authors thank the members of the lab for reviewing our manuscript and offering helpful feedback. We also thank Zada et al. for making a valuable ECoG dataset public.
Supplementary Materials for ”Interpretable Embeddings of Speech Enhance and Explain Brain Encoding Performance of Audio Models”
S1 SSM Feature Extraction
Representations from self-supervised speech models (SSMs) were extracted using the code provided by Antonello et al. (available at https://github.com/HuthLab/encoding-model-scaling-laws). The input audio file, totaling 1800 seconds, was divided into five contiguous 360-second segments before being fed into the SSMs. This segmentation was performed to prevent data leakage between training and test folds in the 5-fold cross-validation scheme.
We used the largest variants of each SSM (Whisper, WavLM, and HuBERT) because prior findings from Antonello et al. (Antonello, Vaidya, and Huth 2023) showed that the encoding performance of the brain by SSMs increased logarithmically with the number of model parameters. The details of each model are shown in Table.S1.
S2 Interpretable auditory features
Full details of the processing procedures for each interpretable feature set are provided below.
Mel-spectrogram
We computed mel-spectrograms from 0–8 kHz with 32 mel bins and further converted them to decibels using librosa (McFee et al. 2015). To create a feature vector for time $t$ s, we concatenated the frames spanning [$t-0.5$, $t$]s, yielding a 672-dimensional vector. Dimensionality was reduced by principal-component analysis (PCA), retaining 95% of the variance.
Gabor bank filter features (GBFB)
Spectro-temporal modulation features were extracted with the Gabor filter-bank implementation of Schädler et al (Schädler, Meyer, and Kollmeier 2012). The 455-dimensional features, computed at 100 Hz, were down-sampled to 4 Hz to align with the neural data; PCA (95% variance) was applied prior to modelling.
Speech presence
Speech probability was estimated with EfficientAT, a CNN-based audio tagger (Schmid, Koutini, and Widmer 2023). To produce a feature scalar for time $t$ s, we fed the model the audio spanning [$t-0.5$, $t$]s.
Phonetic features
Each word in the transcript was converted to its ARPAbet phoneme representations using nltk.corpus.cmudict (Bird and Loper 2004). Word intervals were evenly subdivided among their $N$ phones, assigning on- and offset times to every phoneme. Every phoneme was one-hot encoded over the 39 ARPAbet symbols. To create a feature vector for time $t$ s, the embeddings for every phoneme whose offset time lies in [$t-0.5$, $t$]s were summed, producing a $T\times 39$ matrix.
Semantic features
Following Benara et al. 2024, we obtained 168-dimensional question-answering (QA) semantic embeddings by prompting GPT-4 (gpt-4-0125-preview; OpenAI 2023) with 56 yes/no questions about each word’s meaning (e.g., “Does the word/sentence involve an expression of personal values or beliefs?”) with 3 different context lengths (word-level, 1.5 s, and 3 s). For the 1.5/3-second context length, the model was additionally provided with the words spoken in the 1.5/3 seconds preceding each target word when answering the questions. To create a feature vector for time $t$ s, the embeddings for every word whose offset time lies in [$t-0.5$, $t$]s were summed, producing a $T\times 168$ matrix. Because of how this feature is constructed, its effective context window can extend up to approximately 4.75 seconds—accounting for the 3-second input context, the additional 0.5-second summing window, and the maximum word duration (about 1.25 seconds for words like “approximately”). A full list of questions is provided in Table S2.
Syntactic features
Following Zada et al. 2025, each word was first converted into 96-dimensional embeddings, comprising 50 part-of-speech tags, 45 dependency labels, and a stop/non-stop label using the spaCy (Honnibal et al. 2020) library and its en_core_web_lg model. To create a feature vector for time $t$ s, the embeddings for every word whose offset time lies in [$t-0.5$, $t$]s were summed, producing a $T\times 96$ matrix.
S3 Semantic Question-Answering Features
The list of all questions for generating semantic Question-Answering features is provided in Table.S2. The categories were assigned based on clustering on their cosine similarity. Questions 36-56 were not assigned any category as they were deemed less relevant for neural encoding than the first 35 questions.
Fig. S6 shows the results of category-level analysis, in which questions in QA features were divided into sub-categories (Miscellaneous, Visual, Social, Numeric, Abstract / Opinion) based on the clustering of these questions. The results show that SSMs encode semantic information broadly rather than being limited to any specific semantic sub-category, provided the information is neurally relevant.
S4 Variance Partitioning
Between two feature sets $\mathcal{A}$ and $\mathcal{B}$,
$R^{2}_{\mathcal{A}_{\textbf{unique}}}$ $=R^{2}_{\mathcal{(A,B)}}-R^{2}_{\mathcal{B}}$
$R^{2}_{\mathcal{B}_{\textbf{unique}}}$ $=R^{2}_{\mathcal{(A,B)}}-R^{2}_{\mathcal{A}}$
$R^{2}_{{}_{\textbf{shared}}}$ $=R^{2}_{\mathcal{A}}+R^{2}_{\mathcal{B}}-R^{2}_{\mathcal{(A,B)}}$
where $R^{2}_{\mathcal{A}}$, $R^{2}_{\mathcal{B}}$, and $R^{2}_{\mathcal{(A,B)}}$ denote the variance explained by models using only $\mathcal{A}$, only $\mathcal{B}$, and both $\mathcal{A}$ and $\mathcal{B}$ together, respectively. All models were fitted using banded ridge regression.
S5 Effect of context length on brain prediction performance
Among the context sizes tested, the context size with the highest average correlation was 30 seconds for Whisper and HuBERT, and 5 seconds for WavLM (Fig. S4). This variability highlights the need for caution when interpreting the influence of context length on encoding performance. For instance, Anderson et al. found that EEG encoding performance with Whisper peaked at context lengths of 5–10 seconds and interpreted this as evidence that EEG preferentially reflects the encoding of speech within that temporal window. However, it is essential to distinguish whether such patterns reflect inherent properties of neural processing or are instead artifacts of model-specific biases.
@@ Acknowledgements
Figure S1: Unique variance explained $R^{2}$ by each feature set. The language processing hierarchy can be clearly observed in the anatomical distributions.
Figure S2: Overall Average Pearson’s $r$ across electrodes for different models with the best layer and context size. The difference was found to be statistically significant ($p<0.001$) with paired t-test.
Figure S3: Average Pearson’s $r$ in Heschl’s gyrus and Insula for the interpretable model and early/late layers of SSMs. We assessed statistical significance using paired t-tests followed by FDR correction.
Figure S4: Average Pearson’s $r$ across all electrodes for the different context sizes of SSMs. Left: Whisper, Middle: WavLM, Right: HuBERT
Figure S5: (a) Variance $R^{2}$ of mel-spectrogram [$t-0.5$, $t$]s explained by HuBERT representations over different layers. (b) Average variance explained $R^{2}$ by HuBERT over the 0.2 seconds [$t-0.2$, $t$]s for different frequency bins, plotted over layers. The gain/loss in percentage with respect to the 1st layer is displayed on the middle plot. On the right plot, the gain/loss percentage for the 37th layer is shown. (c, d): Corresponding plots for WavLM, analogous to (a) and (b), respectively.
Figure S6: Unique variance $R^{2}$ explained by 5 categories of semantic features in IFG-Tri relative to non-semantic interpretable features and the joint model with SSM representations.
@@
Table S1: Summary of the self-supervised speech models used in this study.
@@
HuBERT X-large 48 1280 964M Masked prediction of quantised latent units from iterative k-means clustering of MFCCs
WavLM Large 24 1024 317M Same masked-unit objective as HuBERT $+$ utterance-mixing and denoising augmentation
Whisper Large 32 1280 638M Supervised multilingual encoder–decoder ASR and speech-to-text translation
Table S2: List of all questions in question-answering features and their categories. N/A means not assigned any category.
Question Category
1. Does the sentence describe a personal reflection or thought? Abstract / Opinion
2. Does the sentence contain a proper noun? Misc
3. Does the sentence describe a physical action? Visual
@@
4. Does the sentence describe a personal or social interaction that leads to a change or revelation? Social
5. Does the sentence involve the mention of a specific object or item? Visual
6. Does the sentence involve a description of physical environment or setting? Visual
7. Does the sentence describe a relationship between people? Social
8. Does the sentence mention a specific location? Misc
9. Is time mentioned in the input? Numeric
10. Is the sentence abstract rather than concrete? Abstract / Opinion
@@
11. Does the sentence express the narrator’s opinion or judgment about an event or character? Misc
12. Is the input related to a specific industry or profession? Misc
13. Does the sentence include dialogue? Social
14. Does the sentence describe a visual experience or scene? Visual
15. Does the input involve planning or organizing? Misc
16. Does the sentence involve spatial reasoning? Visual
17. Does the sentence involve an expression of personal values or beliefs? Abstract / Opinion
18. Does the sentence contain a negation? Abstract / Opinion
19. Does the sentence describe a sensory experience? Visual
20. Does the sentence include technical or specialized terminology? Misc
21. Does the input contain a number? Numeric
22. Does the sentence contain a cultural reference? Misc
23. Does the text describe a mode of communication? Social
24. Does the input include a comparison or metaphor? Abstract / Opinion
@@
25. Does the sentence express a sense of belonging or connection to a place or community? Misc
26. Does the sentence describe a specific sensation or feeling? Misc
27. Does the text include a planning or decision-making process? Misc
28. Does the sentence include a personal anecdote or story? Misc
@@
29. Does the sentence involve a discussion about personal or social values? Abstract / Opinion
30. Does the text describe a journey? Misc
31. Does the input contain a measurement? Numeric
32. Does the sentence describe a physical sensation? Misc
33. Does the sentence include a direct speech quotation? Social
34. Is the sentence reflective, involving self-analysis or introspection? Abstract / Opinion
35. Does the input describe a specific texture or sensation? Misc
36. Is the input about a discovery or realization? N/A
37. Does the sentence include an account of a miscommunication or misunderstanding? N/A
38. Does the sentence include a specific sound or auditory description? N/A
39. Does the sentence use a unique or unusual word? N/A
40. Does the sentence describe a change in a physical or emotional state? N/A
41. Does the sentence describe a moment of relief or resolution of tension? N/A
42. Does the sentence include a conditional clause? N/A
43. Does the sentence reference a specific time or date? N/A
44. Is the sentence conveying the narrator’s physical movement or action in detail? N/A
45. Is there mention of a city, country, or geographic feature? N/A
46. Does the sentence involve an unexpected incident or accident? N/A
47. Does the sentence involve a recount of a social or community event? N/A
48. Does the sentence express a philosophical or existential query or observation? N/A
49. Does the story involve a personal project or creation? N/A
50. Is the sentence emotionally positive? N/A

### v2

@@ 4 Discussion
In this work, we examined the encoding model performance of speech foundation models (SFMs) for human auditory and language processing. By systematically comparing the largest variants of three state-of-the-art models (Whisper, HuBERT, and WavLM) with six classes of handcrafted acoustically/linguistically interpretable features in their ability to predict ECoG responses, our analysis advances the understanding of how SFMs encode speech information relevant for neural processing. Crucially, this work addresses several previously unresolved aspects of SFM encoding performance, clarifying their representational alignment with human perceptual and linguistic processes.
First, we demonstrated that the vast majority of neural variance captured by speech foundation models (SFMs) is attributable to interpretable features—nearly all in the earliest layers, and still a substantial portion (approximately 80%) in the deeper layers. This finding suggests that much of the observed SFM–brain alignment reflects well-characterized acoustic and linguistic structure. In a parallel line of work, Hadidi et al. (2025) reported similar results in the context of LLMs, showing that most of the neural variance explained by GPT2-XL could be accounted for by simple confounds such as word-rate and positional information. Together, these results indicate that much of the apparent model–brain alignment across modalities can be traced back to straightforward, interpretable features. The unique neural variance explained by SFMs—beyond what is accounted for by interpretable features—is primarily localized to the superior temporal gyrus and inferior frontal gyrus, and increases with model size. Moving forward, it will be important to precisely characterize the nature of this residual variance, for example by applying linear probing methods to the residual SFM representations after regressing out interpretable features.
Second, our analyses revealed a systematic trade-off in how SFMs encode neurally relevant information across layers: the shared neural variance with low-level features (mel-spectrogram, GBFB) peaked in early layers, while it increased monotonically with higher-order features (phonetic, syntactic, semantic) over layers. A closer look at mel-spectrogram encoding showed that Whisper and WavLM progressively lose fine-grained acoustic detail but expand their temporal receptive windows with depth, whereas HuBERT exhibits wide windows from layer 0. These differences likely stem from model architectures: HuBERT’s convolutional relative positional embeddings give tokens long receptive windows from layer 0; WavLM’s additional gating mechanism likely explains its divergence from HuBERT; and Whisper’s use of fixed sinusoidal encodings directly on log-mel spectrograms constrains early receptive windows, forcing them to expand only in later layers.
@@
Third, whether SFMs encode brain-relevant semantics is debated, with evidence supporting (Pasad et al., 2022; Chen et al., 2024; Anderson et al., 2024) and challenging (Choi et al., 2024; Oota et al., 2024). Our method employs an interpretable semantic embedding to directly test for the shared neural variance with SFMs, while controlling for lower-level acoustic and linguistic information. We show that SFM representations significantly reduce the unique variance explained by the semantic embedding, providing strong evidence that they indeed encode brain-relevant semantic information. However, an important nuance remains: SFMs account for only about 40% of the unique variance explained by the semantic features in IFG, with respect to non-semantic features, indicating that their semantic representations remain partial. Recent studies (Moussa et al., 2025; Vattikonda et al., 2025) have shown that fine-tuning SFMs on fMRI data can improve their alignment with language regions of the brain. Thus, while our findings affirm the presence of semantic information in SFM representations, they also suggest room for improvement.
Lastly, we showed that augmenting SFM representations with interpretable features significantly improves brain encoding model performance. This strategy overcomes the limitations of relying on individual SFM layers by integrating complementary information from hand-crafted features while still benefiting from the representational advantages of model scaling. Importantly, we do not rule out the possibility that SFM representations alone could achieve comparable performance. Indeed, alternative approaches have aggregated embeddings from multiple tokens or combined representations across layers, as in prior work (Goldstein et al., 2022). However, such strategies are substantially less efficient. For example, because Whisper processes audio as log-mel spectrograms with positional encodings, representing a 0.5-s segment requires approximately 25 tokens (each covering  20 ms). For Whisper-large, this translates to 25 × 1,280 = 32,000 dimensions. By contrast, our interpretable feature set achieves the same encoding with fewer than 1,500 dimensions (672 for the mel-spectrogram alone), while capturing all of the neural variance explained by early SFM layers. In other words, using interpretable features provides a more compact and interpretable representation of the same information, making them a principled complement to SFM embeddings.
@@
Limitations. 1. Our findings are based on data from nine clinical ECoG participants who listened to a single 30‑min English podcast. While the high temporal resolution of ECoG and broad electrode coverage across brain regions make this dataset particularly well-suited for the present study, replication using larger and more diverse datasets will be important for assessing the generalizability of these results. 2. While we controlled for several lower-level features in our analysis of semantic encoding in SFMs, we cannot entirely rule out the possibility that unaccounted confounds contribute to the shared variance between SFM representations and our semantic features. However, the observed effect of context size provides some evidence against this interpretation, as lower-level features typically operate over shorter integration windows and are therefore less sensitive to extended context. 3. While our interpretable features account for most of the neural variance captured by SFM representations, a statistically significant portion remains unexplained. Identifying the nature of this residual variance is an important direction for future research. 4. While there are different classes of models that can take speech as input, including AudioLLMs, we restricted our analyses to SFMs. The rationale is twofold: first, most current AudioLLMs are modular systems that pair an audio encoder—often Whisper or a self-supervised learning model—with a language model, so their audio representations largely build on the same foundations we study. Second, AudioLLMs are primarily optimized for interactive or task-specific use cases, such as question-answering or dialogue, rather than for producing stable, general-purpose audio features, making them less directly suited for the type of neural encoding comparisons considered in this work. Nonetheless, future work may explore how AudioLLMs, given their rapid development and growing capabilities, could provide complementary insights into neural processing of speech.
Broader Impact Statement
This study relies on electrocorticography (ECoG) data collected from patients with epilepsy who underwent invasive monitoring for clinical purposes. While the dataset we analyze is fully de-identified and publicly available, it is important to acknowledge that these recordings originate from a vulnerable patient population. The use of such data requires careful attention to ethical considerations, including the assurance that appropriate informed consent and clinical oversight were in place during data collection, and that secondary analyses such as ours respect the dignity and autonomy of the participants.
Appendix A Appendix
A.1 Unique Neural Variance by Each Interpretable Feature
@@
Figure S1: Unique neural variance $R^{2}$ explained by each feature with respect to the rest of the interpretable features. The language processing hierarchy can be clearly observed in the anatomical distributions.
A.2 SFM Feature Extraction
Representations from speech foundation models (SFMs) were extracted using the code provided by Antonello et al. (2023) (available at https://github.com/HuthLab/encoding-model-scaling-laws). The input audio file, totaling 1800 seconds, was divided into five contiguous 360-second segments before being fed into the SFMs. This segmentation was performed to prevent data leakage between training and test folds in the 5-fold cross-validation scheme.
A.3 Effect of context length on brain prediction performance
Among the context sizes tested, the context size with the highest average correlation was 30 seconds for Whisper and HuBERT, and 5 seconds for WavLM (Fig.˜S2). This variability highlights the need for caution when interpreting the influence of context length on encoding performance. For instance, Anderson et al. found that EEG encoding performance with Whisper peaked at context lengths of 5–10 seconds and interpreted this as evidence that EEG preferentially reflects the encoding of speech within that temporal window. However, it is essential to distinguish whether such patterns reflect inherent properties of neural processing or are instead artifacts of model-specific biases.
Figure S2: Average correlation between predicted and actual brain responses using embeddings from Whisper, WavLM, and HuBERT across different context lengths. The maximal correlations across layers were obtained with context lengths of 30 s for Whisper, 5 s for WavLM, and 30 s for HuBERT.
A.4 Variance partitioning with high-level features
In Fig.˜2c, the ratio for the phonetic, syntactic, and semantic features was computed using three-way variance partitioning among the SFM representations, speech probability, and the respective feature set (see Fig.˜S3), in order to control for speech presence information embedded within these higher-level features. For mel-spectrogram and Gabor filter bank features, no such control was applied, as they by definition represent the lowest-level confounds.
Figure S3: Diagram of the three-way variance partitioning analysis used to quantify the neural variance shared between higher-order features (phonetic, syntactic, and semantic) and SFM representations. Because these higher-order features also encode information about speech presence, we controlled for it explicitly within the partitioning framework.
Figure S4: The ratio of neural variance shared between different interpretable feature sets and the SFM representations, relative to the total variance explained by each interpretable feature set.
Figure S5: Unique neural variance of WavLM and HuBERT with respect to the combined set of all interpretable features, plotted on a template brain.
Figure S6: (a) Average neural variance $R^{2}$ shared between WavLM representations and mel-spectrogram segments of different temporal windows. (b) Variance $R^{2}$ of the mel-spectrogram [t0.5t-0.5, $t$]s explained by the WavLM representations over different layers. (c, d) Corresponding plots for HuBERT.
@@ A.4 Variance partitioning with high-level features
Table S1: Summary of the speech foundation models used in this study.
@@
HuBERT X-large 48 1280 964M
Masked prediction of quantised latent units
from iterative k-means clustering of MFCCs
WavLM Large 24 1024 317M
Same masked-unit objective as HuBERT
++ utterance-mixing and denoising augmentation
Whisper Large 32 1280 638M
Supervised multilingual encoder–decoder ASR
and speech-to-text translation
Table S2: List of all questions in question-answering features.
Question
1. Does the sentence describe a personal reflection or thought?
2. Does the sentence contain a proper noun?
3. Does the sentence describe a physical action?
@@
4. Does the sentence describe a personal or social interaction that leads to a change or revelation?
5. Does the sentence involve the mention of a specific object or item?
6. Does the sentence involve a description of physical environment or setting?
7. Does the sentence describe a relationship between people?
8. Does the sentence mention a specific location?
9. Is time mentioned in the input?
10. Is the sentence abstract rather than concrete?
@@
11. Does the sentence express the narrator’s opinion or judgment about an event or character?
12. Is the input related to a specific industry or profession?
13. Does the sentence include dialogue?
14. Does the sentence describe a visual experience or scene?
15. Does the input involve planning or organizing?
16. Does the sentence involve spatial reasoning?
17. Does the sentence involve an expression of personal values or beliefs?
18. Does the sentence contain a negation?
19. Does the sentence describe a sensory experience?
20. Does the sentence include technical or specialized terminology?
21. Does the input contain a number?
22. Does the sentence contain a cultural reference?
23. Does the text describe a mode of communication?
24. Does the input include a comparison or metaphor?
@@
25. Does the sentence express a sense of belonging or connection to a place or community?
26. Does the sentence describe a specific sensation or feeling?
27. Does the text include a planning or decision-making process?
28. Does the sentence include a personal anecdote or story?
@@
29. Does the sentence involve a discussion about personal or social values?
30. Does the text describe a journey?
31. Does the input contain a measurement?
32. Does the sentence describe a physical sensation?
33. Does the sentence include a direct speech quotation?
34. Is the sentence reflective, involving self-analysis or introspection?
35. Does the input describe a specific texture or sensation?
36. Is the input about a discovery or realization?
37. Does the sentence include an account of a miscommunication or misunderstanding?
38. Does the sentence include a specific sound or auditory description?
39. Does the sentence use a unique or unusual word?
40. Does the sentence describe a change in a physical or emotional state?
41. Does the sentence describe a moment of relief or resolution of tension?
42. Does the sentence include a conditional clause?
43. Does the sentence reference a specific time or date?
44. Is the sentence conveying the narrator’s physical movement or action in detail?
45. Is there mention of a city, country, or geographic feature?
46. Does the sentence involve an unexpected incident or accident?
47. Does the sentence involve a recount of a social or community event?
48. Does the sentence express a philosophical or existential query or observation?
49. Does the story involve a personal project or creation?
50. Is the sentence emotionally positive?